OpenStudy (anonymous):

My answer is A Select the best possible first step to solving the following system by first eliminating the y variable. 4x + 5y = -8 3x + 2y = 1 (4 points) Answer Multiply the first equation by 2 and multiply the second equation by 5. Multiply the first equation by -2 and multiply the second equation by -5. Multiply the first equation by -2 and multiply the second equation by 5. Multiply the first equation by 8 and the second equation by -8.

OpenStudy (phi):

This question is a bit goofy. If we use your choice A, the 5y in the first equation times 2 gives 2*5y= 10y 5 times the 2y in the 2nd equation gives 5*2y= 10y if we subtract the bottom equation from the top (or vice versa) 10y-10y=0 eliminates the y variable however, if we use choice C: multiply the top by -2 and the bottom by +5 we would get -2*5y= -10y and 5*2y= +10y we can add the 2 equations and -10y+10y =0 , and y is eliminated. what is " best possible first step "? is adding easier than subtracting? I would go with choice C , but this is not really "the best", whatever that means

OpenStudy (anonymous):

I was aiming towards that as wellt thereem out there emou tthe ree mout thereem out there